# Webly-Supervised Image Manipulation Localization via Category-Aware Auto-Annotation

### v1

@@ VIII Conclusion
We propose CAAAv2, a novel category-aware framework that solves the constrained image manipulation localization problem by separately handling Shared Donor Group (SDG) and Shared Probe Group (SPG) image pairs through Difference-Aware Semantic Segmentation and Correlation DINO with frozen DINOv2 backbone, dramatically reducing overfitting while enabling high-quality automatic annotation. With CAAAv2 and a novel metric QES, we auto-annotate manually forged images from the web and harvest a large-scale, diverse, high-quality dataset MIMLv2, comprising 246,212 manually forged images with pixel-level annotations. This dataset considerably addresses the data scarcity issue in IML. Furthermore, we propose Object Jitter, an effective data generation method that leverages authentic web images to improve model performance in challenging cases. To better leverage web-scale supervision, we also introduce Web-IML, a new model for IML. Experiments confirm the strong generalization ability of our method, which achieves unprecedented performance gains and surpasses previous state-of-the-art methods by significant margins. We believe this work establishes a scalable framework for continuously expanding training data as new manually forged images emerge online. Its universal applicability is demonstrated across multiple benchmarks. This research transforms image forensics, shifting reliance from limited, expensive handcrafted datasets to abundant, continuously growing web resources, thus paving the way for more robust and generalizable IML systems.
Figure 12: Qualitative comparison on image manipulation localization.

### v2

@@ IX Conclusion
In this paper, we addressed the critical challenge of data scarcity in image manipulation localization through a scalable web-supervision approach. At the core of our approach is CAAAv2, a novel category-aware and prior-denoising auto-annotation paradigm. This innovative design strategically separates manipulations into Shared Donor Group and Shared Probe Group, effectively converting the challenging forensic task into a prior-denoising task. This dramatically reduces task complexity and mitigates overfitting. Leveraging this automatic annotation engine and our proposed QES metric for quality control, we constructed MIMLv2, a diverse, high-quality dataset of 246,212 manually forged images—over 120× larger than existing handcrafted datasets. We further enhanced this web-scale supervision with Object Jitter, a novel technique for generating challenging training examples from authentic images. To fully capitalize on this vast data, we designed Web-IML, a new model featuring multi-scale perception and self-rectification modules. Extensive experiments demonstrate that our web-supervision method provides unprecedented performance gains across multiple models on a wide array of real-world forgery benchmarks. Additionally, Web-IML establishes a new state-of-the-art, significantly outperforming previous methods and demonstrating remarkable generalization. We believe that this work establishes a scalable framework for continuously expanding training data as new manually forged images emerge online. This research transforms image forensics, shifting reliance from limited, expensive handcrafted datasets to abundant, continuously growing web resources, thus paving the way for more robust image manipulation localization systems.